OpenStudy (anonymous):

Two tangents from an external point P are drawn to a circle and intersect it at A and B. A third tangent meets the circle at T, and the tangents PA and PB at points Q and R, respectively. Find the perimeter of Triangle PQR.

OpenStudy (anonymous):

The problem involves Power of a Point theorem.

OpenStudy (anonymous):

Here is the diagram that I made for the question. This should make things easier to understand. Also, this problem requires the use of the Power of a Point theorem, so keep that in mind.

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@campbell_st @Luis_Rivera @phi

OpenStudy (campbell_st):

let PQ = a and AQ = b well you need to know that tangents from an external point are equal so you know AQ = QT and RB = RT and PA = PB therefore PQ = RQ so the triangle is isosceles then the perimeter is P = 2a + 2b thats about the best I can do off the top of my head.

OpenStudy (anonymous):

I did everything you did, the only thing that worried me was that the Perimeter should be some numbers. Now of course that doesn't have to be the case, but I noted everything you did so I am just trying to make sure that the Perimeter in fact is not a number which you also seem to have concluded. If that's the case, then I guess things should be fine.

OpenStudy (campbell_st):

oopss... here is the better solution since QT = AQ and RT = RB then the perimeter is P = PQ + QT + PR + TR or P = PQ + QA + PR + TB = PA + PB since tangents from an enternal point are equal then P = 2PA that makes more sense... hope you can follow it.

OpenStudy (campbell_st):

oops should read P = PQ + QA + PR + RB = PA + PB sorry for the typo

OpenStudy (campbell_st):

then P = 2PA

OpenStudy (anonymous):

Yeah that makes sense. I now realise the only thing that threw me off is the fact that the question is supposed to involve Power of a Point theorem and it doesn't really have much of an application to this question. That's why I kept trying hard to make some sense out of it but turns that the question really doesn't have much to do with it lol. Anyway ty.

OpenStudy (campbell_st):

I'm not familiar with power of a point... I know tangents from an external point are equal... and the logic works... glad to help

OpenStudy (anonymous):

Yes it does indeed. The whole power of a point theorem threw me off and made me think that it was what I needed to solve this question when it actually was not =,=